OpenStudy (anonymous):

If f(x) is a linear function, f(-2) = -5, and f(5) = -4, find an equation for f(x)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (debbieg):

Well, "linear function" just means that it's an equation for a line. So you can use any line form that you like: y=mx+b or \(y-y_1=m(x-x_1)\). Either way, you need the slope and a point. You are given f(-2) = -5, and f(5) = -4... well, those are just 2 points on the line, right? Remember, this is just function notation. y=f(x). So you are given that y=-5 when x=-2. That means that the point (-2,-5) is one point on the line. Get the other point from f(5) = -4. Now you have two points, so you can find slope, then use the slope and either point to find the equation for the line.